# Interpretable Spatio-Temporal Features Extraction based Industrial Process Modeling and Monitoring by Soft Sensor

### v1

@@ IV Experimental Results
To validate the efficiency of our method, we test the proposed STCIM in three stages. First, analytical experiments based on two generated datasets are used to demonstrate the effectiveness of STCIM. Then, a real-life dataset from a 330MW power plant is utilized as a comparative experiment to further validate the robustness and efficiency of STCIM. At last, we emphasize the importance of position encoding, far topological coupling among variables, and hyper-parameters in the loss function.
@@ IV-A Experiment Setup
The utilized datasets contain two generated datasets and one real-world dataset. The two generated datasets include a reduced rotary speed dataset and a 3-order synchronous generator dataset, which are both important for the equipment of industrial processes. The comparative experiment has a 330MW coal-fired power plant dataset that reflects a complex industrial process with heat and mass transformation. All the datasets have two operation conditions, and each condition has at least 50,000 points with noise. The model is trained in one operation condition and tested in another to highlight the effectiveness of STCIM. The state variables are not used as input to models.
For each STCIM block, we have different parameters. In the spatio-temporal feature extraction layer block, the features are extracted by two 2D-convolution layers with 64 filters, and each kernel is $3\times 3$. The far topological alignment layer has four heads, and each head has three linear transformations, and each linear layer has 256 neurons. The number of interpretable variables is chosen based on the real physical model. For the decoder, each fully connected layer has 256 neurons and GeLU as an activation function. For all experiments, $\gamma_{1}$ and $\gamma_{2}$ in Eq. (10) are set as 1 and 0.5 separately. During the training process, the latent variables and the target output of STCIM will be integrated into the discrete state-space loss function and then collaboratively optimized by the backpropagation algorithm.
Four more deep learning-based soft sensor modeling methods are explored for comparisons, including two different PINN-based models [24, 29], and two CNN-based models [10, 7], since they have demonstrated their out-performance in soft sensor. Table I shows the number of parameters of models. For each experiment, we train the models with a batch size of 64 or 128 using the ‘NAdam’ optimizer with an initialized learning rate of 0.01. The learning rate decays by a factor of 0.1 every 30 epochs. The max epoch is set to 100. The experiments are implemented in Tensorflow using a CPU Intel i7-11800H CPU at 2.3Hz and an NVIDIA T600 GPU.
@@
Models STCIM PhyLSTMs [24] PIML [29] VCACNN [7] MSACNN [10]
#PRM 0.023M 0.428M 0.025M 0.371M 0.082M
@@ IV-B Analytical Experiments
In this section, we use two equipment models with different levels of difficulty to demonstrate the reliability and scalability of STCIM in modeling and condition monitoring. For long-time-series datasets, STCIM can be used not only in complex industrial processes, but also in relatively simple equipment monitoring without over-fitting. The well-designed modules improve the accuracy of the modeling and monitoring.
@@ IV-B1 Rotor Equation-Reduced Rotary Speed
The stability of the rotor is one of the key factors for the stable operation of the micro gas turbine. During the operation of the micro gas turbine, the rotor needs to maintain a certain speed and stability to ensure the normal operation of the machine. If the rotor vibrates or becomes unstable, it will have a negative impact on the operation of the micro gas turbine and may even cause damage to the machine. Therefore, the modeling and monitoring of the rotor need to be fully considered.
The corresponding reduced rotary speed can be calculated by the rotor equation and its transformation, which is given by
$\begin{split}&\frac{\mathrm{d}n}{\mathrm{d}t}=\frac{900}{\pi^{2}Jn}(N_{T}-N_{C}),\\ &\hat{n}=\frac{n}{\sqrt[2]{\frac{T_{1}}{288}}}.\end{split}\vspace{-0.3cm}$ (11)
@@
Table II shows that in the analytical experiment, the STCIM completely outperforms all the other physics-based deep learning models and CNN-based models. The evaluation metrics of STCIM are much smaller than the evaluation metrics of other models, indicating the superiority of STCIM.
TABLE II: Evaluation metrics of the reduced rotary speed
Evaluation STCIM PhyLSTMs [24] PIML [29] VCACNN [7] MSACNN [10]
MAE $\hat{n}$ 0.14 0.79 1.24 0.24 0.85
$n$ 0.16 1.03 1.33 0.26 0.75
MSE $\hat{n}$ 0.03 1.01 2.28 0.10 0.46
$n$ 0.04 1.68 2.57 0.12 0.63
[…]
$n$ 5.73 36.84 41.59 9.32 28.64
@@
The synchronous generator is the core equipment in the power system, responsible for converting mechanical energy into electrical energy. Its stable operation is directly related to the power supply quality and reliability of the entire power system. At the same time, since the synchronous generator can stably output the required voltage by adjusting the excitation current, and the rotor is synchronized with the power grid, it is crucial to protect power equipment and ensure the stability of the power supply.
A 3-order synchronous generator is a practical model that has been widely used in modeling and monitoring. We leverage the classic 3-order synchronous generator as the simulation model, which is given as follows for data generation.
@@
where $i$ is the input current, $u$ is the output voltage, $w$ is the angular velocity and $\hat{E_{q}}$ is the transient internal voltage. Since $\delta$ is not related to outputs, it will be ignored in our case. Other parameters are constants. The generated dataset contains all the generated currents, voltages, angular velocity, and power with 50,000 points for each operation condition.
The Table III shows the evaluation of the state estimation ($w,\hat{E_{q}}$) and output regression ($i_{q},i_{d}$) across five different architectures. The proposed STCIM outperforms the other deep learning methods among most evaluation metrics, especially in the output regression and $w$ estimation. Compared with other models, the performance of $\hat{E_{q}}$ estimation of STCIM is comparable.
TABLE III: Evaluation metrics of the 3-order synchronous generator
Evaluation STCIM PhyLSTMs [24] PIML [29] VCACNN [7] MSACNN [10]
MAE ($10^{-2}$) $i_{q}$ 3.38 3.45 3.57 3.41 3.66
$i_{d}$ 3.63 4.07 4.23 3.78 4.05
$w$ 9.18 16.94 12.51 10.27 11.03
$\hat{E}_{q}$ 1.46 1.45 1.47 1.42 1.33
MSE ($10^{-3}$) $i_{q}$ 1.75 1.79 1.88 1.80 1.99
$i_{d}$ 2.08 2.55 2.73 2.62 2.72
$w$ 13.12 46.30 57.13 36.78 18.05
$\hat{E}_{q}$ 0.32 0.35 0.34 0.35 0.36
[…]
$i_{d}$ 27.03 28.44 29.32 28.01 27.21
$w$ 9.18 16.93 17.29 16.87 11.02
$\hat{E}_{q}$ 1.52 1.52 1.48 1.51 1.49
@@ IV-C Comparative Experiments
This section is used to demonstrate the effectiveness of STCIM in long-time-series datasets of complex industrial processes. The real-life dataset comes from a 330MW coal-fired power plant, which is shown in Figure 5. The processes of coal-fired power plants are multifaceted and complex, including chemical reactions, energy conversion, mass transfer, and phase changes. The utilized 330MW power plant [30] can be given by
[…]
To verify the generalization of STCIM, it is tested under both the same operation condition (Test 1) and a different operation condition (Test 2). Figure 6 shows the outputs and state estimation of the real-life dataset. To reveal the performance of STCIM, 400 points in the test dataset are selected in the figures. Table IV compares the STCIM with four other deep learning models in the dataset. Compared with Figure 6a, the performance of STCIM in Figure 6b is better owing to the obvious fluctuations in state variables and the detailed information from sub-devices. As depicted in the Figures, the changes in output power and the amount of coal are larger, promoting the learning of system dynamic characteristics by STCIM. The test of STCIM under the same operation conditions shown in Figure 6a has good performance, including the forecasting of state variables. Meanwhile, STCIM maintains good generalization ability in prediction under a new operation condition.
Table IV shows the evaluation metrics among the five deep learning models in both Test 1 and Test 2. When the models are tested under the same operation condition, the performance of all the tested models is similar. Some models can outperform STCIM in some specific evaluation metrics because the dataset under the same operation condition has similarity, for example, noise in the output. To pursue better generalization, STCIM pays more attention to consistent spatio-temporal features and meaningful latent variables rather than the noise. In contrast, under another operation condition, the generality of STCIM is much better than that of other models. All evaluation metrics for all power prediction and latent variables prediction by STCIM are smaller than other methods, implying the effectiveness of STCIM.
Figure 5: 330MW coal power plant.
(a) Test 1 of the real-life dataset
(b) Test 2 of the real-life dataset
@@
Figure 6: The outputs and latent variables of comparative experiments
TABLE IV: Evaluation metrics of the real-life dataset
Evaluation STCIM PhyLSTMs [24] PIML [29] VCACNN [7] MSACNN [10]
MAE (Test 1) $y$ 0.45 0.41 0.40 0.46 0.78
$z_{1}$ 0.43 0.76 0.61 0.44 1.07
$z_{2}$ 0.18 0.21 0.69 0.18 0.51
MSE (Test 1) $y$ 0.31 0.29 0.26 0.31 1.01
$z_{1}$ 0.25 0.87 0.59 0.29 1.73
$z_{2}$ 0.05 0.08 1.71 0.06 0.43
[…]
$z_{1}$ 1.20 1.98 1.58 1.12 2.77
$z_{2}$ 0.98 1.12 1.61 0.99 2.66
MAE (Test 2) $y$ 0.46 0.54 0.57 0.46 0.75
$z_{1}$ 0.11 0.56 0.14 0.28 0.68
$z_{2}$ 0.05 0.05 1.38 0.11 0.26
MSE (Test 2) $y$ 0.49 0.50 0.67 1.08 0.94
$z_{1}$ 0.02 0.07 0.03 0.05 0.07
$z_{2}$ 0.03 0.04 0.07 0.16 0.09
[…]
$z_{1}$ 0.35 0.65 0.35 0.44 0.55
$z_{2}$ 0.33 0.34 0.54 0.67 0.61
@@ IV-D Ablation Experiments
In this section, we discuss the importance of spatio-temporal consistency and the hyper-parameters of the physics-informed interpretable latent space block in STCIM. The real-life dataset is utilized in ablation experiments. The setup is the same as in comparative experiments.
IV-D1 The importance of spatio-temporal consistency
In this experiment, we emphasize the importance of spatio-temporal consistency by comparing models with and without position encoding using a real-life dataset, which is shown in Figure 7. Compared with Figure 6, there is little difference in output $y$ forecasting under the same operation condition. However, the predicted state variables without adaptive position encoding are not as accurate as the model with it. A similar situation appears in Test 2. The volatility of all the predicted variables, including latent variables and output, is greater than that of the model with position encoding, especially in actual fuel forecasting. The bias between field data and predicted actual fuel in Figure 7b is much greater than that in Figure 6b.
Table V shows the evaluation metrics of experiments in the real-life dataset without adaptive position encoding. Compared with Table IV, Test 1 with or without position encoding has similar performance. Nevertheless, the generalization of the model without adaptive position encoding in Test 2 is worse compared to the model performance in Table IV. For example, the MSE in $y$ forecasting without adaptive position encoding almost triples compared with the model with adaptive position encoding. The improvement of generalization benefits from adaptive position encoding. By position encoding, the features are endowed with additional information, which improves the expressiveness of STCIM.
(a) Test 1 for the real-life dataset without position encoding
(b) Test 2 for the real-life dataset without position encoding
Figure 7: The outputs and latent variables of experiments in a real-life dataset without position encoding.
TABLE V: Evaluation metrics of experiments in real-life dataset without position encoding
[…]

### v2

@@ IV Experimental Results
To validate the efficiency of our method, we test the proposed STCIM in three stages. First, analytical experiments based on two generated datasets are used to demonstrate the effectiveness of STCIM. Then, a real-life dataset from a 330MW power plant is utilized as a comparative experiment to further validate the robustness and efficiency of STCIM. The scalability of STCIM is tested using a dataset from 1000MW ultra-supercritical once-through boiler-turbine unit. At last, we emphasize the importance of position encoding, far topological coupling among variables, and hyper-parameters in the loss function.
@@ IV-A Experiment Setup
The two generated datasets include a rotor equation-reduced rotary speed dataset and a 3-order synchronous generator dataset, which are both important for the equipment of thermal power plant. The comparative experiment has 330MW and 1000MW coal-fired power plant datasets that reflects a complex industrial process with heat and mass transformation. All the datasets have two operation conditions, and each condition has at least 50,000 points with noise. The model is trained in one operation condition and tested in another to highlight the effectiveness of STCIM. The state variables are not used as input to models.
For each STCIM block, we have different parameters. In the local spatio-temporal feature extraction layer block, the features are extracted by two 2D-convolution layers with 8 filters, and each kernel is $3\times 3$. The far topological alignment layer has four heads, and each head has three linear transformations, and each linear layer has 32 neurons. The number of interpretable variables is chosen based on the real physical model. For the decoder, each fully connected layer has 32 neurons and GeLU as an activation function. For all experiments, $\gamma_{1}$ and $\gamma_{2}$ in Eq. (9) are set as 1 and 0.5 separately. During the training process, the latent variables and the target output of STCIM will be integrated into the discrete state-space loss function and then collaboratively optimized by the backpropagation algorithm.
The STCIM is composed of two sub-blocks with computationally intensive expressions: the local spatio-temporal feature extraction block and the far topological alignment block. For the attribute size of input, $X(t)\in R^{H\times W\times C}$, the computation will intensify in the local spatio-temporal feature extraction block by $\mathcal{O}(K^{2}\times C\times M\times H\times W)$ factor, where the K is the kernel size, M is the number of output channels of the block. Similarly, the computation complexity of the far topological alignment block is $\mathcal{O}(D\times B\times H\times(W^{2}+H\times W+W\times m))$, where D is the total number of attention layers and feedforward layers, B is the batch size, $m$ is the hidden dimension of the MLP. Since the local spatio-temporal feature extraction block only contains the two 2D-convolution layers, the computational complexity of STCIM is mainly affected by far topological alignment block which contains attention and feedforward layers. The attention mechanism’s $\mathcal{O}(H^{2}\times W)$ term may be significant for long sequences, while the $\mathcal{O}(W^{2})$ term dominates for large model dimensions.
Six more deep learning-based soft sensor modeling methods are explored for comparisons, including PhyLSTM [26], PIML [31], MSACNN [15], VCACNN [17], MIAN [11], and TimeVAE [32]. The six methods include the PINN-based interpretable models and the CNN-based, attention-based, and VAE-based spatiotemporal features extraction models. They have demonstrated their out-performance in different soft sensor of large-scale industrial processes. Table I shows the number of parameters of models. For each experiment, we train the models with a batch size of 64 or 128 using the ‘NAdam’ optimizer with an initialized learning rate of 0.01. The learning rate decays by a factor of 0.1 every 30 epochs. The max epoch is set to 100. The experiments are implemented in Tensorflow using a CPU Intel i7-11800H CPU at 2.3Hz and an NVIDIA T600 GPU.
@@
Models STCIM PhyLSTMs [26] PIML [31] VCACNN [17]
#PRM 0.023M 0.428M 0.025M 0.371M
Models MSACNN [15] MIAN [11] TimeVAE [32] -
#PRM 0.082M 0.012M 0.302M -
@@ IV-B Analytical Experiments
In this section, we use two widely used equipment models in thermal power plants to demonstrate the reliability and scalability of STCIM in modeling and condition monitoring. For the evaluation, we use MAE, MSE and MAPE as the indicators.
TABLE II: Evaluation metrics of the 3-order synchronous generator
Evaluation STCIM PhyLSTMs [26] PIML [31] VCACNN [17] MSACNN [15] MIAN [11] TimeVAE [32]
MAE ($10^{-2}$) $i_{q}$ 3.38 3.45 3.57 3.41 3.66 3.52 3.69
$i_{d}$ 3.63 4.07 4.23 3.78 4.05 3.95 4.07
$w$ 9.18 16.94 12.51 10.27 11.03 10.54 14.12
$\hat{E}_{q}$ 1.46 1.45 1.47 1.42 1.33 1.40 1.42
MSE ($10^{-3}$) $i_{q}$ 1.75 1.79 1.88 1.80 1.99 1.83 1.94
$i_{d}$ 2.08 2.55 2.73 2.62 2.72 2.65 2.71
$w$ 13.12 46.30 57.13 36.78 18.05 25.45 39.78
$\hat{E}_{q}$ 0.32 0.35 0.34 0.35 0.36 0.36 0.36
MAPE $i_{q}$ 4.15% 4.24% 4.38% 4.26% 4.47% 4.34% 4.51%
$i_{d}$ 27.03% 28.44% 29.32% 28.01% 27.21% 27.68% 27.46%
$w$ 9.18% 16.93% 17.29% 16.87% 11.02% 15.48% 13.98%
$\hat{E}_{q}$ 1.52% 1.52% 1.48% 1.51% 1.49% 1.50% 1.49%
@@ IV-B1 Rotor Equation-Reduced Rotary Speed
The stability of the rotor is one of the key factors for the stable operation of thermal power plants. During the operation, the rotor needs to maintain a certain speed and stability to ensure the normal operation of the machine and avoid the damage to the machine. The reduced rotary speed can be given by
$\begin{split}&\frac{\mathrm{d}n}{\mathrm{d}t}=\frac{900}{\pi^{2}Jn}(N_{T}-N_{C}),\\ &\hat{n}=\frac{n}{\sqrt[2]{\frac{T_{1}}{288}}}\end{split}$
@@
Fig 3 shows that in a simple equipment monitoring, the STCIM completely outperforms all other models, including the physics-based and feature extraction-based models. First, the conventional fully connected network is not enough to extract the necessary features, resulting in the worst performance of PIML. Then, the time-related features can improve the forecasting precession, which is reflected in PhyLSTM and TimeVAE. Due to the limited variables and the fast characteristics, the local or the global spatio-temporal features in the rotor equation are similar, which makes the performances of VCACNN, MSACNN, and MIAN similar. However, in this experiment, the underlying physical laws in PhyLSTM and PIML do not improve the accuracy of state variable estimation, which may be due to the simplicity of the models.
Figure 3: Evaluation metrics of rotor equation reduced rotary speed.
@@
The synchronous generator is the core equipment in the thermal power system, responsible for converting mechanical energy into electrical energy. Its stable operation is directly related to the reliability and safety of the entire power system. We use the classic 3-order synchronous generator as the simulation model, which is given as follows for data generation.
@@
where $i$ is the input current, $u$ is the output voltage, $w$ is the angular velocity, and $\hat{E_{q}}$ is the transient internal voltage. Since $\delta$ is not related to outputs, it will be ignored in our case. Other parameters are constants. The generated dataset contains all the generated currents, voltages, angular velocity and power with 50,000 points for each operation condition.
Table II shows the evaluation of the state estimation ($w,\hat{E_{q}}$) and the output regression ($i_{q},i_{d}$). The proposed STCIM outperforms the other deep learning methods in most evaluation metrics, especially in the output regression and $w$ estimation. Compared to other models, the performance of $\hat{E_{q}}$ estimation of STCIM is competitive as well. The bias in the latent variable $\hat{E_{q}}$ may be caused by the hyper-parameter $\gamma_{2}$ of loss function. Considering the good performance of $w$ estimation and that the dataset of $3^{rd}$ order synchronous generator is generated, the hyper-parameter $\gamma_{2}$ can be set larger for better performance. Similar to fig 3, due to the local or global spatio-temporal feature extraction, VCACNN, MSACNN, and MIAN have better performance than PhyLSTM, PIML, and TimeVAE. As the complexity of the model increases, accurate feature extraction becomes increasingly important. Another interesting thing is that the regression accuracy of state estimation ($w,\hat{E_{q}}$) of PhyLSTM and PIML is becoming competitive, which is different from Fig 3 . We need to emphasize that the regression of state estimation ($w,\hat{E_{q}}$) of PhyLSTM and PIML is unsupervised, since we add the discrete state-space loss in the final loss function like STCIM. They are fundamentally different from VCACNN, MSACNN, MIAN, and TimeVAE.
TABLE III: Evaluation metrics of the real-life dataset
Evaluation STCIM PhyLSTMs [26] PIML [31] VCACNN [17] MSACNN [15] MIAN [11] TimeVAE [32]
MAE (Test 1) $y$ 0.45 0.41 0.40 0.46 0.78 0.51 0.58
$z_{1}$ 0.43 0.76 0.61 0.44 1.07 0.67 0.92
$z_{2}$ 0.18 0.21 0.69 0.18 0.51 0.35 0.49
MSE (Test 1) $y$ 0.31 0.29 0.26 0.31 1.01 0.52 0.65
$z_{1}$ 0.25 0.87 0.59 0.29 1.73 0.46 0.98
$z_{2}$ 0.05 0.08 1.71 0.06 0.43 0.26 0.37
MAPE (Test 1) $y$ 0.37% 0.26% 0.25% 0.45% 0.78% 0.52% 0.76%
$z_{1}$ 1.20% 1.98% 1.58% 1.12% 2.77% 1.69% 2.35%
$z_{2}$ 0.98% 1.12% 1.61% 0.99% 2.66% 1.32% 2.03%
MAE (Test 2) $y$ 0.46 0.54 0.57 0.46 0.75 0.59 0.66
$z_{1}$ 0.11 0.56 0.14 0.28 0.68 0.32 0.58
$z_{2}$ 0.05 0.05 1.38 0.11 0.26 0.18 0.22
MSE (Test 2) $y$ 0.49 0.50 0.67 1.08 0.94 0.92 0.87
$z_{1}$ 0.02 0.07 0.03 0.05 0.07 0.07 0.06
$z_{2}$ 0.03 0.04 0.07 0.16 0.09 0.12 0.10
MAPE (Test 2) $y$ 0.14% 0.15% 0.21% 0.48% 0.23% 0.37% 0.29%
$z_{1}$ 0.35% 0.65% 0.35% 0.44% 0.55% 0.48% 0.52%
$z_{2}$ 0.33% 0.34% 0.54% 0.67% 0.61% 0.65% 0.62%
@@ IV-C Comparative Experiments
This section is used to demonstrate the effectiveness of STCIM in a real thermal power plant. We collect the dataset from a real 330MW thermal power plant, which follows the introduction in Section II. In the dataset, there are two operating conditions and 13 features, and each feature has 96,000 points with noise. The generalization of STCIM is tested under both the same operation condition (Test 1) and a different operation condition (Test 2).
IV-C1 The comparisons of different methods
Table III summarizes the evaluation metrics of all models in both Test 1 and Test 2. At the same operating point, the performance of the models varies, which might be caused by the size of models, the different captured features. For models that only focus on local features (MSACNN), the performance is almost the worst when predicting the output and the state variables at the same time. For long time-series data, the models related time-related features (PhyLSTM and TimeVAE) outperform MSACNN, indicating the importance of long-term characteristics in the data. In PIML, time as input and domain decomposition help it achieve good performance, especially the output forecasting. Models with spatio-temporal feature achieve the best performance among all models, including STCIM, VCACNN, and MIAN. Benefiting from the built-in physics-informed interpretable latent space loss function, STCIM has more accurate state variable monitoring without labels. However, in another operating point, the generalization of STCIM is much better than that of other models. An interesting thing is that the generalization of the physic-informed models (PhyLSTM and PIML) in this experiment is better than other feature extraction-based models, especially the state variables forecasting. It indicates that when the thermal power plant follows unified physical laws, embedding the physical information is equivalent to extracting the intrinsic spatio-temporal features, without complex network architecture, which helps the model to generalize more easily.
IV-C2 The visualization of output and interpretable latent variables
Fig. 4 shows the output forecasting and state estimation (interpretable latent variables) of the real thermal power plant dataset using STCIM. To reveal the performance of STCIM, 400 points in the test dataset are selected in the figures, including Test 1 and Test 2. Although the performance of STCIM in output forecasting in Test 1 is not the best, it still closely follows the actual value. Compared with Fig. 4a, the performance of STCIM in Fig. 4b is better, owing to the obvious fluctuations in state variables, the detailed information from sub-devices,and the physic-informed loss function. Therefore, as well as the spatio-temporal feature consistency, the explicit interpretable latent space along with the physic-informed loss function helps STCIM maintain a good generalization ability in prediction and monitoring under a new operating point.
(a) Test 1 of the real thermal power plant dataset
(b) Test 2 of the real thermal power plant dataset
@@ IV-C2 The visualization of output and interpretable latent variables
Figure 4: The output forecasting and state variables monitoring of comparative experiments
IV-C3 The scalability of STCIM: 1000MW Thermal Power Plant
To demonstrate that STCIM can be applied to other thermal power plants, we use a real dataset from a 1000MW ultra-supercritical once-through boiler-turbine unit [33]. The inputs are separately the fuel signal (kg/s), total water flow (kg/s), turbine valve opening (%). The latent variables ($z_{1},z_{2}$) are separately the steam-water separator pressure (MPa) and the steam-water separator enthalpy (kJ/kg). The outputs ($y_{1},y_{2}$)are the main steam pressure (MPa) and the output power (MW). Each features has 96,000 points with noise.
Fig 5 summarizes the evaluation metrics of all models when tested under a different operation condition. STCIM outperforms other models in this experiment. Since the process of the 1000MW ultra-supercritical once-through boiler-turbine unit is more complex than that of the 330MW thermal power plant, the differences between the models are more significant. The integrated physical information loss function forces the models (STCIM, PhyLSTM, and PIML) to satisfy the first principles, which improves the generalization ability of the models, especially in steam-water separator enthalpy forecasting. Instead of extracting spatio-temporal features from a single operation condition, the additional loss and the adopted spatio-temporal consistency can provide more potential physical information to ensure the safety of thermal power plants.
Figure 5: Evaluation metrics of 1000MW thermal power plant.
@@ IV-D Ablation Experiments
In this section, we discuss the importance of spatio-temporal consistency and the hyper-parameters of the physics-informed interpretable latent space block in STCIM. The real-life dataset of 330MW thermal power plant is utilized in ablation experiments. The setup is the same as in comparative experiments.
IV-D1 The importance of adaptive position encoding
Table IV shows the evaluation metrics of Test 1 and Test 2 without adaptive position encoding. Compared with Table III, there is a slight decrease in the output forecasting and variable estimation, indicating the effectiveness of spatio-temporal consistency. Whereas, the difference is not remarkable. In contrast, the generalization of STCIM without adaptive position encoding in Test 2 is worse compared to the performance of STCIM in Table III. For example, the MSE in $y$ forecasting without adaptive position encoding almost triples compared with STCIM with adaptive position encoding. The improvement of generalization benefits from adaptive position encoding. By position encoding, the features are endowed with additional information, which improves the expressiveness of STCIM.
TABLE IV: Evaluation metrics of experiments in real-life dataset without position encoding
Datasets MSE MAE MAPE